# xxxxx, xxxxx, xxxx

### v3

@@ 4 Relevant Causal Identities
The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language. That language is dispositional [bennettmaruyama2022a], shaped by natural selection, and so a weaker policy is likely to be a fit policy. WPO will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
@@ 4.1 Ascribing Intent to Other Objects
The distinction between “intervention” and not is misleading. Passive observation is just bearing witness to an intervention by something other than one’s self. The question is not “is this an intervention" but “by whom was this intervention made?".
@@
Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself. However, now we can see that Bob must also construct an identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$. For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [bennett2023d], to satisfy the incentive precondition for causal identity.
@@ 4.3 Realising Lower Order States And Higher Order Meta Representations
Each policy an organism $\mathfrak{o}$ learns implies $\mathfrak{v}$-tasks. A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [sep-peirce-semiotics, bennett2023d] of sign, referent and interpretant. We formalise this as a “protosymbol” [^11] Proto because it is something more primitive than a symbol as conceived of by Peirce. system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$ (see definitions 6 and 7).
Tasks exist in a “generational hierarchy”. They are not mutually exclusive. Higher level tasks are more general, and have fewer policies because only very weak policies could complete them. Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations. Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [goertzel2006]. Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
@@ 5 Multi-Layered Self-Organisation
As the vocabulary and capacity for WPO scale, a greater variety of concepts can be learned. Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [bennett2023c]. This lets us frame the construction of embodied selves in developmental [ciaunica2019] and evolutionary terms.
@@ 5.1 The First Order Self
A first order self (1ST henceforth, formalised in definition 13) allows an organism discern the consequences of its actions. This serves as the locus of self related information processing and experience [bennett2023c], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. 1ST order selves amount to reafference [merker2007, barron2016], observed in mammals and insects.
@@
A 1ST order self may require centralisation and a “solid brain” with a persistent structure [sole2019] (an individual human), as opposed to a “liquid brain” (e.g. an ant colony or a population of humans).
@@ 5.2 The Second Order Selves
Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [bennett2023d]). This is called a second order self (2ND henceforth). We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [grice1957, grice1969]. Grice argued that communication is about the inference of intent. If person $\mathfrak{a}$ and $\mathfrak{b}$ are talking, then the meaning $m_{\mathfrak{a}}$ of what $\mathfrak{a}$ says is whatever $\mathfrak{a}$ intends $\mathfrak{b}$ understand. The meaning $m_{\mathfrak{b}}$ that $\mathfrak{b}$ understands is whatever $\mathfrak{b}$ thinks $\mathfrak{a}$ wants $\mathfrak{b}$ to think. $\mathfrak{b}$ has understood what $\mathfrak{a}$ means if $m_{\mathfrak{b}}$ approximates $m_{\mathfrak{a}}$. This can happen only if $\mathfrak{a}$ can predict with reasonable accuracy what $\mathfrak{b}$ thinks $\mathfrak{a}$ thinks, and $\mathfrak{b}$ can predict what $\mathfrak{a}$ thinks $\mathfrak{b}$ will think upon hearing an utterance. In other words, both $\mathfrak{a}$ and $\mathfrak{b}$ must have 2ND order selves that are good approximations. Yes, there are other aspects to communication.
@@
Where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report. One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^13] This is ‘meta-self-reflexive consciousness’ as some have described it [morin2006]..
@@
More formally using the notation given in the appendix [bennett2024appendices] (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour. Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part. Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part. What is important here is that each organism’s intent is to some extent inferred by the other, changing the sorts of policies that are fit. For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [alexander2022] [^14] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [bennett2023d], and transient relations, information asymmetry and other factors can make deceit a more attractive option.. Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [bennett2023d]. To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent. The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
@@ 5.3 The Third Order Selves
We can scale preconditions indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [morin2006]. This is the awareness that one is self-aware. If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD. Formally a 3RD order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$. It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
@@ 6 The What and Why of Consciousness
It follows that every policy learned in this way must classify a quality. Hence every such policy is a local state. A causal identity imbued with disposition by valence. There would be a policy for one’s act of smelling coffee. For perceiving one’s friend. There would be something different it is like to interact with a hostile version of that very same friend. The 1ST order self accompanies everything an organism does. It has a quality, so the 1ST order self is “what it is like” to be that organism. Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [nagel1974]. An organism does not make a decision to interpret information, it just reacts. What we call subjective experience is the process of learning and enacting a lattice of causal identities, and there is something that “has” all of those subjective experiences once there is a 1ST order self that is part of them all. This is where phenomenal consciousness begins, with experience as learning and exploration.
@@
Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [merker2005, merker2007, barron2016]. We agree that reafference is key, but we provide a very different explanation of why and how. Their work presents biological evidence for subjective experience in organisms with reafference. In contrast, we derive the 1ST order self from first principles and explain why and how it is a classifies “what it is like” to be a particular organism. The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises. Hence these are complementary positions.
@@ 7 Unifying Lower and Higher Order Theories of Consciousness
The psychophysical principle of causality takes us from attraction to and repulsion from physical states, to higher order meta representations. This produces different orders of self. We suggest that the full richness of human subjective experience depends upon these different orders of self [^15] Perhaps the full richness of “hard” consciousness [boltuc2012] depends upon the interaction of selves, perhaps in planning [shanahan2012]..
@@
Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour. Cells are a material with agency [ball2023], and human intelligence is the high level goal directed behaviour of a swarm of cells [fields2020, levin2021, levin2024]. When we embody human abstractions (e.g. arithmetic) in silico (e.g. the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [bennett2024c]. We call the information embodied in the computer “representational” because it means something to us. However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [bennett2024b]. This suggests there is no such thing as representational contents. They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature: the causes of valence. This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive!
@@ 8 From Rocks to Einstein: The Hierarchy of Being
Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist [heylighen2023b] in a stable environment.
@@
Stage two introduces learning. To learn an organism must store, classify and order historical examples by valence. However there is not something it is like to be stage two, because there is no locus of “self”. A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora. Even Tripedalia cystophora was recently shown to be capable of associative learning [bielecki2023]. An entirely distributed control system can “learn”. Likewise, stage two is exemplified by nematodes [rankin2022, willett2018] with a centralised nervous system and some ability to adapt with experience. However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities [barron2016]. When starved C. elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [barron2016, luersen2014, artyukhin2015], whereas something like a bee or an ant can recall and navigate to previously discovered food [wehner2013, seeley1995page295, oades1978].
@@
This is where phenomenal consciousness begins, with a 1ST order self. In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have [merker2005, merker2007, barron2016]. They identified a housefly as a good example of where subjective experience may begin, and we concur. We also hold this is where an organism might be said to have intent. Intuitively, the policy that motivated behaviour is the intent of that behaviour [^16] In the same way declarative and imperative programs are equivalent [howard1980].. For example, eating tends to involve the intent of satisfying hunger.
A stage three organism can feel simple things like hunger, but cannot cannot conceive of itself from another’s perspective. Subsequently it cannot communicate in the Gricean sense [grice1969, bennett2023d], or conceive of its own death, or experience shame.
@@
Stage four is the 2ND order self, and this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense. The ability of ravens to intentionally deceive [bugnyar2002] suggests they are at least stage four. Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$. $\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$. It seems likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived. For example, a cat anticipates its prey will flee when it is observed, and hides.
@@
A 3RD order self facilitates complex planning, particularly of a social nature. A 3RD order self is a 2ND order self for one’s 2ND order self. Humans appear to possess this, because we are aware that we are aware. Altruistic behaviour observed in Australian magpies [potvin2022] suggests they may also have 3RD order selves.

### v4

@@ 4 Relevant Causal Identities
The optimal choice of policy for adaptability is the weakest because it is implied by more statements in an embodied language. That language is dispositional [61], shaped by natural selection, and so a weaker policy is likely to be a fit policy. WPO will correctly identify cause and effect where relevant, simplifying the entire environment into objects and properties relevant to the organism’s motivations.
@@ 4.1 Ascribing Intent to Other Objects
The distinction between “intervention” and not is misleading. Passive observation is just bearing witness to an intervention by something other than one’s self. The question is not “is this an intervention” but “by whom was this intervention made?”.
@@
Bob’s causal identity for himself $c_{b}\subset b.\ int-obs$ only represents the intervention by himself. However, now we can see that Bob must also construct an identity $c_{l}$ for Larry, where the $c_{l}\subset l.\ int-obs$. For an organism $\mathfrak{a}$ with an embodied language $L_{\mathfrak{v}_{\mathfrak{a}}}$ to construct a causal identity for an object $\mathfrak{b}$, it must first be the case that $\mathfrak{a}$ is affected by $\mathfrak{b}$ [50], to satisfy the incentive precondition for causal identity.
@@ 4.3 Realising Lower Order States And Higher Order Meta Representations
Each policy an organism $\mathfrak{o}$ learns implies $\mathfrak{v}$-tasks. A $\mathfrak{v}$-task is a triadic relation between inputs, outputs and policies which resembles Peircean semiosis [62, 50] of sign, referent and interpretant. We formalise this as a “protosymbol” [^11] Proto because it is something more primitive than a symbol as conceived of by Peirce. system $\mathfrak{s}_{\mathfrak{o}}$ for the organism $\mathfrak{o}$ (see definitions 6 and 7).
Tasks exist in a “generational hierarchy”. They are not mutually exclusive. Higher level tasks are more general, and have fewer policies because only very weak policies could complete them. Related as they are in a lattice, protosymbols are analogous lower order states and higher order meta representations. Some have framed consciousness as a problem of moving from unary, to dyadic, to triadic relations [63]. Similarly, we have gone from unary states to dyadic programs, to triadic tasks and protosymbols.
@@ 5 Multi-Layered Self-Organisation
As the vocabulary and capacity for WPO scale, a greater variety of concepts can be learned. Progressively higher orders of ‘causal-identity’ for one’s self related information processing are constructed [34]. This lets us frame the construction of embodied selves in developmental [64] and evolutionary terms.
@@ 5.1 The First Order Self
A first order self (1ST henceforth, formalised in definition 13) allows an organism discern the consequences of its actions. This serves as the locus of self related information processing and experience [34], allowing the organism to plan complex interactions and maintain a consistent “self” that is part of the present interaction. 1ST order selves amount to reafference [65, 66], observed in mammals and insects.
@@
A 1ST order self may require centralisation and a “solid brain” with a persistent structure [67] (an individual human), as opposed to a “liquid brain” (e.g. an ant colony or a population of humans).
@@ 5.2 The Second Order Selves
Survival may demand organism $\mathfrak{a}$ infer $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s interventions (to see one’s self as if through another’s eyes [50]). This is called a second order self (2ND henceforth). We argue that if access conscious contents are available for communication in the human sense, then they must be communicable in the Gricean sense [68, 69]. Grice argued that communication is about the inference of intent. If person $\mathfrak{a}$ and $\mathfrak{b}$ are talking, then the meaning $m_{\mathfrak{a}}$ of what $\mathfrak{a}$ says is whatever $\mathfrak{a}$ intends $\mathfrak{b}$ understand. The meaning $m_{\mathfrak{b}}$ that $\mathfrak{b}$ understands is whatever $\mathfrak{b}$ thinks $\mathfrak{a}$ wants $\mathfrak{b}$ to think. $\mathfrak{b}$ has understood what $\mathfrak{a}$ means if $m_{\mathfrak{b}}$ approximates $m_{\mathfrak{a}}$. This can happen only if $\mathfrak{a}$ can predict with reasonable accuracy what $\mathfrak{b}$ thinks $\mathfrak{a}$ thinks, and $\mathfrak{b}$ can predict what $\mathfrak{a}$ thinks $\mathfrak{b}$ will think upon hearing an utterance. In other words, both $\mathfrak{a}$ and $\mathfrak{b}$ must have 2ND order selves that are good approximations. Yes, there are other aspects to communication.
@@
Where a 1ST order self might allow one to observe a cat and form plans regarding causal interactions with the cat, a 2ND order self would allow one to be consciously aware of the cat for the purpose of reasoning and report. One can know of the cat, and that another organism knows of the cat, but a 2ND order self is insufficient to be aware that one is aware of the cat [^13] This is ‘meta-self-reflexive consciousness’ as some have described it [70]..
@@
More formally using the notation given in the appendix [35] (see quick reference guide in section 2), assume $\mathfrak{a}$ and $\mathfrak{b}$ are organisms that evolved to accurately predict one another’s behaviour. Assume $\mathfrak{a}$ constructs a causal identity $c^{\mathfrak{b}}_{\mathfrak{a}}$ to predict $\mathfrak{b}$ given input $i_{\mathfrak{a}}\in I_{\mu_{\mathfrak{a}}}$, of which a second order self $c^{\mathfrak{ba}}_{\mathfrak{a}}$ is part. Likewise, $\mathfrak{b}$ constructs $c^{\mathfrak{a}}_{\mathfrak{b}}$ to predict $\mathfrak{a}$ given input $i_{\mathfrak{b}}\in I_{\mu_{\mathfrak{b}}}$, of which $c^{\mathfrak{ab}}_{\mathfrak{b}}$ is part. What is important here is that each organism’s intent is to some extent inferred by the other, changing the sorts of policies that are fit. For example, second order self means each knows the other can anticipate manipulation, which means the optimal policy will often be to have rather than feign intent that aligns to some extent with the other party’s desires, to co-operate [71] [^14] Depending upon circumstances, for example organisms may co-operate in some circumstances but not others [50], and transient relations, information asymmetry and other factors can make deceit a more attractive option.. Repeated interaction creates an iterated prisoner’s dilemma, incentivising co-operation and signals that both parties interpret similarly (the beginnings of language) [50]. To communicate in Gricean terms, $\mathfrak{a}$ must intend to convey meaning $m_{\mathfrak{a}}$, and $\mathfrak{b}$ must recognise this intent. The incentive precondition explains why $\mathfrak{a}$ would form such intent (co-operation is often advantageous), while how may be understood as follows:
@@ 5.3 The Third Order Selves
We can scale preconditions indefinitely, however for the purposes of this paper we will stop at 3RD order selves, because this appears to be the highest level of human consciousness, known as Meta self-awareness [70]. This is the awareness that one is self-aware. If self-awareness stems from 2ND order selves, then it follows that meta self-awareness requires 3RD. Formally a 3RD order self for $\mathfrak{a}$ reflecting off $c^{\mathfrak{ba}}_{\mathfrak{a}}$ lets $\mathfrak{b}$ is $c^{\mathfrak{baba}}_{\mathfrak{a}}$. It is $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$’s prediction of $\mathfrak{b}$’s prediction of $\mathfrak{a}$.
@@ 6 The What and Why of Consciousness
It follows that every policy learned in this way must classify a quality. Hence every such policy is a local state. A causal identity imbued with disposition by valence. There would be a policy for one’s act of smelling coffee. For perceiving one’s friend. There would be something different it is like to interact with a hostile version of that very same friend. The 1ST order self accompanies everything an organism does. It has a quality, so the 1ST order self is “what it is like” to be that organism. Put another way, Nagel’s question of “what it is like to be” a particular organism could be answered if one could somehow have that organism’s first order self [1]. An organism does not make a decision to interpret information, it just reacts. What we call subjective experience is the process of learning and enacting a lattice of causal identities, and there is something that “has” all of those subjective experiences once there is a 1ST order self that is part of them all. This is where phenomenal consciousness begins, with experience as learning and exploration.
@@
Our arguments align with those of Merker, Barron and Klein, who have linked subjective experience to reafference [72, 65, 66]. We agree that reafference is key, but we provide a very different explanation of why and how. Their work presents biological evidence for subjective experience in organisms with reafference. In contrast, we derive the 1ST order self from first principles and explain why and how it is a classifies “what it is like” to be a particular organism. The 1ST order self also happens to be equivalent to reafference, so we arrive at the same conclusion as Merker from very different, mathematical premises. Hence these are complementary positions.
@@ 7 Unifying Lower and Higher Order Theories of Consciousness
The psychophysical principle of causality takes us from attraction to and repulsion from physical states, to higher order meta representations. This produces different orders of self. We suggest that the full richness of human subjective experience depends upon these different orders of self [^15] Perhaps the full richness of “hard” consciousness [73] depends upon the interaction of selves, perhaps in planning [74]..
@@
Far from suggesting there is no such thing as qualia, this suggests instead that there is no such thing as purely representational content in anything but the tools we construct and our interpretations of their behaviour. Cells are a material with agency [75], and human intelligence is the high level goal directed behaviour of a swarm of cells [76, 77, 23]. When we embody human abstractions (e.g. arithmetic) in silico (e.g. the x86 instruction set), we disconnect the high level goal directed behaviour from the low level behaviour (and the motivating valence) that gave rise to it [38]. We call the information embodied in the computer “representational” because it means something to us. However when we embody our abstractions in silico we disconnect them from the valence that motivated their construction [36]. This suggests there is no such thing as representational contents. They are a fiction we have invented because we struggle to reduce our own abstractions to their basic nature: the causes of valence. This ‘dissolves’ the hard problem by going a level down to the fundamental drive: stay alive!
@@ 8 From Rocks to Einstein: The Hierarchy of Being
Stage one refers to adaptations hard-wired by natural selection, allowing complexity to persist [78] in a stable environment.
@@
Stage two introduces learning. To learn an organism must store, classify and order historical examples by valence. However there is not something it is like to be stage two, because there is no locus of “self”. A biological example of such a decentralised nervous system is the cubozoan box jellyfish Tripedalia cystophora. Even Tripedalia cystophora was recently shown to be capable of associative learning [79]. An entirely distributed control system can “learn”. Likewise, stage two is exemplified by nematodes [80, 81] with a centralised nervous system and some ability to adapt with experience. However, the absence of a “self” limits causal reasoning, which as others have already pointed out must limit spatial, navigational abilities [66]. When starved C. elegans exhibit “increased locomotion and dispersal in a random, rather than directed, search” [66, 82, 83], whereas something like a bee or an ant can recall and navigate to previously discovered food [84, 85, 86].
@@
This is where phenomenal consciousness begins, with a 1ST order self. In biological terms this implies reafference, which others have argued is the key to subjective experience, albeit for different reasons than what we have [72, 65, 66]. They identified a housefly as a good example of where subjective experience may begin, and we concur. We also hold this is where an organism might be said to have intent. Intuitively, the policy that motivated behaviour is the intent of that behaviour [^16] In the same way declarative and imperative programs are equivalent [37].. For example, eating tends to involve the intent of satisfying hunger.
A stage three organism can feel simple things like hunger, but cannot cannot conceive of itself from another’s perspective. Subsequently it cannot communicate in the Gricean sense [69, 50], or conceive of its own death, or experience shame.
@@
Stage four is the 2ND order self, and this is where we hold access consciousness begins because it is where information is available for report in the Gricean sense. The ability of ravens to intentionally deceive [87] suggests they are at least stage four. Raven $\mathfrak{a}$, aware that it is being observed by raven $\mathfrak{b}$, will act as if it is hiding food in one location to mislead $\mathfrak{b}$, but will then move the food in another location unobserved by $\mathfrak{b}$. $\mathfrak{a}$ seems to predict not just the intent of $\mathfrak{b}$ (to steal the food), but $\mathfrak{b}$’s perception of $\mathfrak{a}$. It seems likely that dogs and cats have second order selves, as they must hunt reasonably intelligent animals and must anticipate how their actions are perceived. For example, a cat anticipates its prey will flee when it is observed, and hides.
@@
A 3RD order self facilitates complex planning, particularly of a social nature. A 3RD order self is a 2ND order self for one’s 2ND order self. Humans appear to possess this, because we are aware that we are aware. Altruistic behaviour observed in Australian magpies [88] suggests they may also have 3RD order selves.